OpenStudy (anonymous):

help.... Which of the following is the slope between the points (-2, 4) and (2, -2)?

OpenStudy (sweetburger):

lol i copy and pasted the wrong formula

OpenStudy (anonymous):

I already graphed it

OpenStudy (sweetburger):

use slope formula Slope Formula

jimthompson5910 (jim_thompson5910):

hint: use the slope formula m = (y2 - y1)/(x2 - x1) in this case x1 = -2 y1 = 4 x2 = 2 y2 = -2

OpenStudy (sweetburger):

y2-y1/x2-x1

OpenStudy (anonymous):

can't i graph it then to rise over run?

jimthompson5910 (jim_thompson5910):

Graphing works as well. The slope formula is a way to do it without the graph. But if you're more comfortable with the graph, then go ahead and do it that way.

OpenStudy (anonymous):

kay thanks

jimthompson5910 (jim_thompson5910):

np, tell me what you get

OpenStudy (anonymous):

I got -2/3

OpenStudy (anonymous):

I got -6/2 but simplified it to -2/3

jimthompson5910 (jim_thompson5910):

You have the x and y mixed, up. You should have gotten this m = (y2-y1)/(x2-x1) m = (-2-4)/(2-(-2)) m = (-2-4)/(2+2) m = -6/4 m = -3/2

OpenStudy (anonymous):

that is what i got but i wrote it wrong sorry

jimthompson5910 (jim_thompson5910):

oh hmm..not sure how you got -6/2 when you should have gotten -6/4 that's ok